# Gaussian Variational Inference with Non-Gaussian Factors for State Estimation: A UWB Localization Case Study

### v1

@@ V-C Experimental Results
The experimental evaluation follows the same process and measurement models used in the simulation study presented in Section V-A. Specifically, the process dynamics remain as defined in (25), while the UWB measurement model follows (26), with parameters identified from the empirical error distributions aggregated across all ten trajectories shown in Figure 4. Each estimator was initialized using dead-reckoned gyroscope and wheel odometry measurements prior to evaluation across all cluttered trajectories. The results, summarized in Figure 5, demonstrate that ESGVI provides improved positional accuracy and comparable consistency and orientation accuracy relative to the asymmetric Cauchy MAP estimator. Notably, unlike in the simulation study, the GMM-based estimator performs less effectively in the experimental setting. This degradation is likely due to varying proportions of NLOS measurements across trajectories, which hinders the GMM’s ability to generalize, in contrast to the Monte Carlo simulations where NLOS measurements occurred at a fixed rate of 25%. From Figure 3, it is evident that estimator performance deteriorates when the vehicle traverses NLOS regions caused by the metal, while other obstacles appear to have a less pronounced effect on the trajectory. On the other hand, both the asymmetric Cauchy MAP and ESGVI estimators exhibit comparable orientation RMSE, consistent with the findings in [esgvi_paper], which demonstrated that the performance advantages of the vector-valued ESGVI algorithm over MAP methods are primarily in translational accuracy rather than orientation estimation. This trend appears to extend to the MLG case as well. Regarding consistency, all estimators tend to be slightly overconfident, with the asymmetric Cauchy MAP and ESGVI exhibiting the most similar behavior. This overconfidence is expected because the metric is computed separately for each independent trajectory, and the proportion of NLOS measurements varies from one trajectory to another, making it a challenging dataset.
@@ V-D Limitations
Although ESGVI demonstrates improved accuracy, it is computationally more expensive than MAP-based methods, often an order of magnitude slower in practice. As noted in [esgvi_paper], ESGVI has the same theoretical asymptotic complexity as MAP, since both require a lower–diagonal–upper decomposition of the information matrix. However, the cubature-based computation of marginal expectations increases the constant factor, resulting in higher runtimes. In practice, this is further exacerbated by the lack of speed-oriented optimizations and the use of a simple backtracking line search.
While warm-starting ESGVI from a MAP solution improves convergence, the current implementation remains restricted to offline use and is therefore unsuitable for real-time applications. For a quantitative runtime evaluation, please refer to the relevant section of [esgvi_paper]. Nevertheless, ESGVI remains valuable in scenarios where robustness and improved estimation performance outweigh computational cost, and accelerating ESGVI remains an open direction for future work.
@@ VI Conclusion
In this letter, a more flexible and MLG compatible ESGVI method was presented. Building on the work of [esgvi_paper], it was shown how ESGVI can naturally incorporate skewed measurement noise models while supporting states defined on MLGs. The method was evaluated in an outlier-rich UWB localization experiment, demonstrating greater accuracy and comparable consistency relative to on-manifold robust MAP solutions using GMMs or asymmetric Cauchy losses to handle outliers. Although experimentally ESGVI was more computationally intensive than MAP, these results highlight its potential as a robust and flexible framework for state estimation in challenging environments.

### v2

@@ V-C Experimental Results
The experimental evaluation follows the same process and measurement models used in the simulation study presented in Section V-A. Specifically, the process dynamics remain as defined in (25), while the UWB measurement model follows (26), with parameters identified from the empirical error distributions aggregated across all ten trajectories shown in Figure 4. Each estimator was initialized using dead-reckoned gyroscope and wheel odometry measurements prior to evaluation across all cluttered trajectories. The results, summarized in Figure 5, demonstrate that ESGVI provides improved positional accuracy and comparable consistency and orientation accuracy relative to the asymmetric Cauchy MAP estimator. Notably, unlike in the simulation study, the GMM-based estimator performs less effectively in the experimental setting. This degradation is likely due to varying proportions of NLOS measurements across trajectories, which hinders the GMM’s ability to generalize, in contrast to the Monte Carlo simulations where NLOS measurements occurred at a fixed rate of 25%. From Figure 3, it is evident that estimator performance deteriorates when the vehicle traverses NLOS regions caused by the metal, while other obstacles appear to have a less pronounced effect on the trajectory. On the other hand, both the asymmetric Cauchy MAP and ESGVI estimators exhibit comparable orientation RMSE, consistent with the findings in [2], which demonstrated that the performance advantages of the vector-valued ESGVI algorithm over MAP methods are primarily in translational accuracy rather than orientation estimation. This trend appears to extend to the MLG case as well. Regarding consistency, all estimators tend to be slightly overconfident, with the asymmetric Cauchy MAP and ESGVI exhibiting the most similar behavior. This overconfidence is expected because the metric is computed separately for each independent trajectory, and the proportion of NLOS measurements varies from one trajectory to another, making it a challenging dataset.
@@ V-D Limitations
Although ESGVI demonstrates improved accuracy, it is computationally more expensive than MAP-based methods, often an order of magnitude slower in practice. As noted in [2], ESGVI has the same theoretical asymptotic complexity as MAP, since both require a lower–diagonal–upper decomposition of the information matrix. However, the cubature-based computation of marginal expectations increases the constant factor, resulting in higher runtimes. In practice, this is further exacerbated by the lack of speed-oriented optimizations and the use of a simple backtracking line search.
While warm-starting ESGVI from a MAP solution improves convergence, the current implementation remains restricted to offline use and is therefore unsuitable for real-time applications. For a quantitative runtime evaluation, please refer to the relevant section of [2]. Nevertheless, ESGVI remains valuable in scenarios where robustness and improved estimation performance outweigh computational cost, and accelerating ESGVI remains an open direction for future work.
@@ VI Conclusion
In this letter, a more flexible and MLG compatible ESGVI method was presented. Building on the work of [2], it was shown how ESGVI can naturally incorporate skewed measurement noise models while supporting states defined on MLGs. The method was evaluated in an outlier-rich UWB localization experiment, demonstrating greater accuracy and comparable consistency relative to on-manifold robust MAP solutions using GMMs or asymmetric Cauchy losses to handle outliers. Although experimentally ESGVI was more computationally intensive than MAP, these results highlight its potential as a robust and flexible framework for state estimation in challenging environments.